OpenStudy (anonymous):

let 0 = theta If csc0 = 2 and cos < 0, then sec0+tan0 / sec0 - tan0 = ?

OpenStudy (anonymous):

 sec0=(2/sqrt3) tan0=(1/sqrt3) answer=(3/sqrt3)/(1/sqrt3)=3/1=3